# Effects of retardation on many-body superradiance in chiral waveguide QED

### v1

@@ Abstract
where $v>0$ is the group velocity of the guided modes, $\Gamma$ is the single-atom decay rate into the waveguide, and $x_{n}$ denotes the position of the $n$th atom. We label the atoms according to their position, with $x_{n}<x_{n+1}$.
@@
While these scalings were already proposed in Ref. [4], we are now in a position to confirm them numerically. In Fig. 1c, we show the dependence of $r_{\mathrm{eff}}$ and $t_{\mathrm{eff}}$ on $\tau$, and find that, indeed, at sufficiently small $\Gamma\tau$, the scalings (4) are observed. At larger delays, however, we find that for sufficiently large initial inversion angles $\theta_{0}$, the scalings diverge from the prediction (4). This can be understood by noting that the scalings $r_{n}^{\mathrm{pk}}\sim n\Gamma$ and $t_{n}^{\mathrm{pk}}\sim 1/(n\Gamma)$ hold for $n\gg 1$, however at large $\Gamma\tau$ the effective cooperative system size $N_{\mathrm{eff}}$ becomes small enough for these scalings to break down. This effect is more pronounced for larger $\theta_{0}$, where the asymptotic scalings are approached more slowly.
@@
For partially-inverted initial states, the emission dynamics are also captured by a mean field theory (MFT) description in the limit of small delay times ($\Gamma\tau\ll 1$), where the atomic array can be treated as a continuous medium [3, 6, 4, 52, 7, 8, 1]. We express $\langle\sigma_{n}\rangle_{t}=\frac{1}{2}\sin\theta(n,\Gamma t)$ and $\langle\sigma_{n}^{z}\rangle_{t}=-\frac{1}{2}\cos\theta(n,\Gamma t)$, in terms of a continuous bivariate function $\theta(x,t)$ with dimensionless arguments obeying (see SM [38] and Refs. [55, 56])
@@
Beyond capturing departures from mean-field predictions, our numerics also allow us to study the fully-inverted initial state |$\ket{\psi_{0}}=\ket{e}^{\otimes N}$ conventionally considered in Dicke superradiance, which is adynamical under Eq. (5), since the decay of this state is initialised by vacuum fluctuations not accounted for by MFT [47]. In Figs. 2a,b we see that the plateau effect described above for partially-inverted states manifests analogously in this case. However, the scaling of $t_{\mathrm{eff}}$ displays a significant divergence from the predicted scaling. This can be understood by noting that the correct $n$-dependence of $t^{\mathrm{pk}}_{n}$ in the retardation-free case is actually given by $t^{\mathrm{pk}}_{n}=c_{1}\log(c_{2}n)/n$, with some constants $c_{1},c_{2}$ [59]. Indeed, we find that the scaling $t_{\mathrm{eff}}=c_{3}(\tau/\Gamma)^{1/2}\log(c_{4}/(\Gamma\tau))$ implied by this is a much better fit (see Fig. 2c).

### v2

@@ Abstract
where $v>0$ is the group velocity of the guided modes, $\Gamma$ is the single-atom decay rate into the waveguide, and $x_{n}$ denotes the position of the $n$th atom. We label the atoms according to their position, with $x_{n}<x_{n+1}$. We assume here perfect coupling into the waveguide but demonstrate the robustness of our results in the presence of additional decay into unguided modes in the End Matter.
@@
While these scalings were already proposed in Ref. [4], we are now in a position to confirm them numerically. In Fig. 1c, we show the scalings of $r_{\mathrm{eff}}$ and $t_{\mathrm{eff}}$ with $\tau$, and find that, indeed, at sufficiently small $\Gamma\tau$, the scalings (4) are observed. At larger delays, however, we find that for sufficiently large initial inversion angles $\theta_{0}$, the scalings diverge from the predicted scalings. This can be understood by noting that the scalings $r_{n}^{\mathrm{pk}}\sim n\Gamma$ and $t_{n}^{\mathrm{pk}}\sim 1/(n\Gamma)$ hold for $n\gg 1$, however at large $\Gamma\tau$ the effective cooperative system size $N_{\mathrm{eff}}$ becomes small enough for these scalings to break down. This effect is more pronounced for larger $\theta_{0}$, where the asymptotic scalings are approached more slowly.
@@
For partially-inverted initial states, the emission dynamics are also captured by a mean field theory (MFT) description in the limit of small delay times ($\Gamma\tau\ll 1$), where the atomic array can be treated as a continuous medium [3, 6, 4, 52, 7, 8, 1, 55, 56]. We express $\langle\sigma_{n}\rangle_{t}=\frac{1}{2}\sin\theta(n,\Gamma t)$ and $\langle\sigma_{n}^{z}\rangle_{t}=-\frac{1}{2}\cos\theta(n,\Gamma t)$, in terms of a continuous bivariate function $\theta(x,t)$ with dimensionless arguments obeying (see SM [38] and Refs. [57, 58])
@@
Beyond capturing departures from mean-field predictions, our numerics also allow us to study the fully-inverted initial state |$\ket{\psi_{0}}=\ket{e}^{\otimes N}$ conventionally considered in Dicke superradiance, which is adynamical under Eq. (5), since the decay of this state is initialised by vacuum fluctuations not accounted for by MFT [47]. In Figs. 2a,b we see that the plateau effect described above for partially-inverted states manifests analogously in this case. However, the scaling of $t_{\mathrm{eff}}$ displays a significant divergence from the predicted scaling. This can be understood by noting that the correct $n$-dependence of $t^{\mathrm{pk}}_{n}$ in the retardation-free case is actually given by $t^{\mathrm{pk}}_{n}=c_{1}\log(c_{2}n)/(n\Gamma)$, with some constants $c_{1},c_{2}$ [61]. Indeed, we find that the scaling $t_{\mathrm{eff}}=c_{3}(\tau/\Gamma)^{1/2}\log(c_{4}/(\Gamma\tau))$ implied by this is a much better fit (see Fig. 2c).
@@ Acknowledgements
END MATTER
In realistic settings, the waveguide-coupled atoms are typically also subject to independent decay into unguided free-space modes, which competes with the correlated emission into the waveguide. We account for this effect by supplementing the unitary dynamics under the Hamiltonian (1) with a local dissipation term $\gamma\sum_{n}\mathcal{D}[\sigma_{n}]$ with $\mathcal{D}[x]~{}\bm{\cdot}=x\bm{\cdot}x^{\dagger}-\{x^{\dagger}x,\bm{\cdot}\}/2$. At the level of the exact master equation (2), this amounts to the modification $\mathcal{L}_{n}\hat{\rho}(t)\to\mathcal{L}_{n}\hat{\rho}(t)+\gamma\mathcal{D}[\sigma_{n}]\hat{\rho}(t)$. Here, we discuss the effect of this additional loss mechanism on the retardation-induced peak emission plateau.
Specifically, we consider local emission rates into the waveguide, which we define as $r_{n}^{\mathrm{wg}}(t)=r_{n}(t)-\gamma\langle\sigma_{n}^{\dagger}\sigma_{n}\rangle_{t}$. In the absence of retardation, superradiant decay of the $n$-th atom (i.e. a peak local emission rate $r_{n}^{\mathrm{wg,pk}}>\Gamma$ into the waveguide) can only occur for $n\geq\gamma/\Gamma+2$ [11], as we confirm in Fig. 4a. Beyond this threshhold value, the asymptotic Dicke scaling $r_{n}^{\mathrm{wg,pk}}\sim n\Gamma$ is recovered for sufficiently large $n$.
When $\tau>0$, we observe a more intricate interplay between collective and independent emission, owing to the additional loss of excitations into unguided modes during the propagation time of the photons in the waveguide. As shown in Fig. 4a, at sufficiently small time delays the burst criterion from the retardation-free regime still holds and at sufficiently large $n$ the plateau effect can be observed. However, the plateau emission rate $r^{\mathrm{wg}}_{\mathrm{eff}}$ varies with $\gamma$ and, most notably, there exists a critical free-space decay rate $\gamma_{\mathrm{max}}$ that depends on the retardation time $\tau$, such that for larger $\gamma>\gamma_{\mathrm{max}}$, the burst vanishes entirely for all values of $n$ and $r_{n}^{\mathrm{wg,pk}}=\Gamma$ (see Fig. 4b).
It is worth noting that the emission rates $r_{n}^{\mathrm{wg}}(t)$ into the waveguide are not typically directly measurable in an experimental context. The local field intensity $I_{n}(t)$ in the waveguide (defined as in Eq. (7) of the main text, but with $r_{n}\to r_{n}^{\mathrm{wg}}$), represents a more experimentally accessible quantity. In Fig. 4c, we show that the qualitative signatures of the plateau effect in the peak intensity $I_{n}^{\mathrm{pk}}$ are robust even in the presence of comparatively strong free-space decay.
Figure 4: (a) Peak emission rate into the waveguide for $\tau=0$ (top) and $\Gamma\tau=5\times 10^{-4}$ (bottom). The retardation-free superradiance criterion from Ref. [11] is indicated in white. (b) Plateau emission rate $r_{\mathrm{eff}}^{\mathrm{wg}}$ as a function of free-space decay rate $\gamma$ for various time delays. (c) Local field intensity along the waveguide for $\Gamma\tau=10^{-3}$, both without free-space decay and with $\gamma/\Gamma\in\{1,2,5,10\}$, as indicated by the arrow.